### v1

Abstract
Advances in deep learning for molecular generation show promise in accelerating drug discovery. Bayesian Flow Networks (BFNs) have recently shown impressive performance across diverse chemical tasks, with their success often ascribed to the paradigm of modeling in a low-variance parameter space. However, the Bayesian inference-based strategy imposes limitations on designing more flexible distribution transformation pathways, making it challenging to adapt to diverse data distributions and varied task requirements. Furthermore, the potential for simpler, more efficient parameter-space-based models is unexplored. To address this, we propose a novel Parameter Interpolation Flow model (named PIF) with detailed theoretical foundation, training, and inference procedures. We then develop MolPIF for structure-based drug design, demonstrating its superior performance across diverse metrics compared to baselines. This work validates the effectiveness of parameter-space-based generative modeling paradigm for molecules and offers new perspectives for model design.
keywords:
molecule generation, drug discovery, parameter interpolation flow
1 Introduction
Computer-aided drug design (CADD) has emerged as a pivotal strategy in modern drug discovery and development, extensively employed across various stages—from target identification and validation to lead discovery, optimization, and preclinical evaluation [1, 2, 3]. Among CADD methodologies, structure-based drug design (SBDD) has demonstrated remarkable effectiveness in identifying lead compounds [4]. Nevertheless, CADD-based approaches still face significant challenges, particularly in efficiently leveraging large, imbalanced datasets from public repositories and exhaustively exploring the vast chemical space encompassing diverse ligand conformations and binding poses. Recent advances in Artificial Intelligence Generated Content (AIGC) techniques—successfully applied in text, image, and audio processing [5, 6, 7]—have inspired the development of AI-driven molecular generation frameworks [8, 9, 10, 11]. These innovative approaches enable the extraction of intricate chemical insights from crystallographic data, facilitating the exploration of uncharted chemical territories [12, 13, 14]. By generating novel molecular structures, they significantly expand the accessible chemical space, offering unprecedented opportunities for drug discovery.
Three-dimensional (3D) generative models have demonstrated significant promise in SBDD [15, 16, 17]. They excel by incorporating protein pocket constraints during molecule generation, transferring knowledge across targets to enhance novelty, and enabling end-to-end automation without separate docking steps [18]. For SBDD tasks in 3D space, current mainstream approaches primarily utilize autoregressive or diffusion-based generative frameworks [19, 20, 21]. However, autoregressive models are inherently optimized for data with sequential dependencies and often suffer from mode collapse when applied to unordered data (e.g., molecular structures or images) [22, 23]. Conversely, although diffusion models [24, 25, 26] and flow matching (FM) techniques [27] have demonstrated significant progress in continuous variable modeling, they face unique challenges when applied to molecular generation, which is inherently a multimodal task. The atomic feature space comprises heterogeneous physical quantities spanning distinct data modalities, including discrete variables (e.g., atom types), integer variables (e.g., formal charges), and continuous variables (e.g., spatial coordinates) [28]. Diffusion models and FMs—which operate directly in sample space—encounter optimization challenges when handling discrete data due to the non-differentiable nature of discrete noise perturbations.
To address these limitations, MolCRAFT [29] employs Bayesian Flow Networks (BFNs) [30, 31] to perform SBDD in a fully continuous parameter space. By replacing conventional posterior Bayesian updates after sampling with Bayesian flow distributions during the sampling process, MolCRAFT mitigates excessive noise introduction and effectively resolves the modeling challenges associated with hybrid continuous-discrete spaces in SBDD. Notably, BFNs are increasingly being adopted as foundational frameworks for generative modeling across diverse chemical tasks [32, 33, 34], consistently yielding robust results. However, the Bayesian inference-based strategy restricts the design of more flexible distribution transformation paths, hindering the model’s capacity to generalize across diverse data distributions and meet varied task requirements. By contrast, such flexibility serves as the critical advantage that sets flow-matching models apart from other paradigms.
Therefore, to overcome the respective limitations of flow matching and BFNs, we integrate their core concepts and introduce a novel and versatile parameter interpolation flow (PIF) model, as illustrated in Fig.  1. Our framework treats the complete data distribution as a superposition of Dirac distributions, each corresponding to an individual data point, and trains the model to learn the underlying distributional parameters. Specifically, PIF employs a probabilistic framework that interpolates between Dirac distributions and a chosen prior (e.g., Gaussian or Dirichlet distributions), thereby defining a continuous trajectory toward the prior. During training, the model learns to predict the parameters of intermediate distributions along this trajectory, conditioned on a randomly sampled timestep $t$, and is optimized via Kullback-Leibler (KL) divergence. At inference time, the process begins from the prior and iteratively refines predictions over predefined steps, gradually converging toward the target data distribution. Compared with conventional generative models, PIF facilitates smooth transformations in the parameter space of distributions, rendering it highly versatile for both continuous and discrete data domains. Furthermore, PIF offers flexible prior selection, enabling adaptation to diverse tasks without requiring complex closed-form derivations.
Molecular Parameter Interpolation Flow (MolPIF) extends the PIF framework to molecular generation by learning the parameter spaces associated with atomic coordinates (modeled as Gaussian distributions) and atomic types (modeled as Dirichlet distributions). During training, We incorporate a geometry-enhanced learning strategy, inspired by prior work [35], to provide atomic-level contextual information of ligands to the model. This approach involves randomly masking subsets of atoms during training, enabling the model to dynamically optimize arbitrary atomic arrangements within a given molecular structure. As a result, MolPIF achieves superior performance in overall quality of the generated molecules.
Empirical evaluations conducted on the CrossDocked2020 dataset [36] demonstrate that MolPIF has comprehensive advantages across four key dimensions: (1) Superior de novo generation capability, producing molecules with superior binding properties and chemical validity; (2) Accurate geometric reproduction of molecular structural distributions, including rings, bond lengths, and bond angles; (3) Flexible adaptation to different prior distributions, where we systematically compared and analyzed the advantages and limitations of using Gaussian versus Laplace distributions as priors for modeling atomic coordinates; (4) Effective lead optimization, demonstrating robust performance in enhancing drug candidate properties.
Our main contributions can be summarized as follows:
  • We introduce Parameter Interpolation Flow (PIF), an efficient generative framework that operates directly in parameter space while offering flexible prior distribution selection. The proposed method combines simplicity with practical effectiveness, for which we provide rigorous theoretical foundations and detailed implementations of both training and inference procedures.
  • By applying PIF to molecular generation, we develop MolPIF, which achieves superior performance across multiple evaluation metrics compared to baseline methods.
  • We highlight the potential of parameter-space-based generative modeling paradigm for SBDD, providing a foundation for future methodological developments.
2 Results and discussion
2.1 An overview of the framework of PIF
Parameter Interpolation Flow (PIF) is a flow-based generative model that operates in the parameter space of probability distributions. The core idea is briefly described as follows (for details, see Methods). Unlike conventional flow models that transform samples directly, PIF constructs an interpolation path between the parameters of a prior distribution $\boldsymbol{\theta}_{\text{prior}}$ and a target data-driven distribution $\boldsymbol{\theta}(\textbf{x}_{\text{data}})$. Specifically, it learns a time-dependent parameter trajectory $\boldsymbol{\theta}_{t}=f(t)\boldsymbol{\theta}(\textbf{x}_{\text{data}})+(1-f(t))\boldsymbol{\theta}_{\text{prior}}$, where $f(t)$ is a monotonic function ensuring $\boldsymbol{\theta}_{0}=\boldsymbol{\theta}_{\text{prior}}$ and $\boldsymbol{\theta}_{1}=\boldsymbol{\theta}(\textbf{x}_{\text{data}})$.
During training, PIF optimizes the model $\boldsymbol{\Phi}$ to predict $\boldsymbol{\theta}(\textbf{x}_{\text{data}})$ from samples drawn at intermediate $\boldsymbol{\theta}_{t}$. The prediction accuracy is evaluated via the KL divergence between the predicted distribution $p(\textbf{x}\mid\hat{\boldsymbol{\theta}}_{t+\Delta t})$ and the true interpolated distribution $p(\textbf{x}\mid\boldsymbol{\theta}_{t+\Delta t})$ (Fig. 1a). At inference, PIF generates samples through an iterative refinement process: starting from $\boldsymbol{\theta}_{\text{prior}}$, the model progressively updates parameters until converging to $\boldsymbol{\theta}(\textbf{x}_{\text{data}})$.
We present a toy dataset experiment in the Appendix A, which demonstrates the superior generative capability of our approach in accurately modeling complex data distributions. These results indicate that PIF maintains competitive generative performance compared to Denoising Diffusion Probabilistic Model (DDPM) [25], FM, Straight-Line Diffusion Model (SLDM) [37], and BFN, revealing its significant potential for broader applications across various domains.
Figure 1: Overview of the MolPIF framework. a, Training procedure of the PIF. At timestep $t$, a Dirac distribution is constructed from a variable x, and its parameters are interpolated with the prior to obtain $\boldsymbol{\theta}_{t}$. The neural network predicts Dirac parameters $\hat{\boldsymbol{\theta}}_{t}$ from the previous output and prior interpolation. Training minimizes the KL divergence between distributions parameterized by $\boldsymbol{\theta}_{t}$ and $\hat{\boldsymbol{\theta}}_{t}$, with samples from $p(\textbf{x}\mid\boldsymbol{\theta}_{t})$ as network input. b, Inference process of MolPIF. Prior distributions for atomic coordinates x and types v are sampled iteratively. The SE(3)-equivariant neural network refines distribution parameters by interpolating predicted Dirac parameters with priors at each timestep. At $t=1$, coordinates/types are sampled directly from the predicted Dirac distributions, and the molecule is assembled via OpenBabel. c, Geometry-enhanced learning strategy employed during MolPIF training. A subset of ligand atoms (probability $p_{m}$) is masked (probability $p_{am}$), excluded from interpolation and fixed as context.
2.2 MolPIF: the molecular generation framework based on PIF
Molecular Parameter Interpolation Flow (MolPIF) adapts the PIF framework to generate 3D molecules within protein pockets by modeling atomic coordinates as Gaussian distributions and atom types via Dirichlet distributions (for details, see Methods). The framework interpolates between prior parameters and target data distributions using a monotonic function $f(t)=1-\gamma^{t}$, with losses for coordinates and types combined through weights $\lambda_{\textbf{x}}$ and $\lambda_{\textbf{v}}$. Conditional generation is enabled by fixing substructures as concatenated inputs, ensuring the model preserves specified fragments while generating complementary atoms.
Fig. 1b illustrates the inference process of MolPIF, which initiates from prior distributions and iteratively refines the distributions parameters until the final timestep T. The generated molecule is then assembled using OpenBabel by sampling from the distribution $p(\hat{\boldsymbol{\theta}}_{T})$.
Additionally, this work introduces a geometry-enhanced training strategy (Fig. 1c), where a subset of atoms is fixed as contextual anchors to assist model training. It synthesizes principles from Masked Autoencoders (MAE) [38] and inpainting methods [39] that are well-established in computer vision. During training, we apply two noise addition strategies: (1) comprehensive atomic noise addition, where noise is uniformly applied to all ligand atoms via interpolation, and (2) random mask-based noise addition, where atoms are masked with probability $P_{am}$, and noise is added only to unmasked atoms. These strategies are weighted by probabilities $P_{f}$ (comprehensive noise) and $P_{m}$ (mask-based noise), with $P_{f}+P_{m}=1$. The random masking acts as an implicit data augmentation mechanism, enhancing the model’s understanding of ligand-specific atomic relationships.
By operating directly on distribution parameters, MolPIF efficiently enforces geometric and chemical constraints critical for protein-ligand interactions.
2.3 Datasets, baselines and metrics
We employed the CrossDocked dataset [36] for both training and evaluation purposes. The original dataset comprises 22.5 million protein-ligand complexes. Following RMSD-based filtering and a 30% sequence identity split as implemented by AR [40], the processed dataset consists of 99,900 training pairs and 100 test proteins. For evaluation purposes, we randomly sampled 100 molecules for each test protein to ensure comprehensive assessment.
For baselines in this study, five SOTA models, AR [40], Pocket2Mol [41], TargetDiff [42], DecompDiff [43], MolCRAFT [29], were used. These approaches represent diverse computational strategies for molecular modeling. AR adopts the perspective of molecular atomic density grids, assigning atomic probabilities to each voxel and generating molecular structures atom by atom using a Markov Chain Monte Carlo (MCMC) [44] method. Pocket2Mol utilizes an auto-regressive scheme to generate continuous 3D atomic positions, furthermore, it incorporates bond generation to produce more realistic molecular structures. TargetDiff is based on both continuous and discrete diffusion denoising probabilistic models to enable the simultaneous generation of entire molecules. DecompDiff leverages virtual point searching for arm and scaffold clustering as prior knowledge, and integrates bond diffusion and validity guidance. MolCRAFT employs the BFNs [30] to address the continuous-discrete gap in modeling.
The performance assessment incorporates three distinct categories of evaluation metrics: (a) Binding Affinity. AutoDock Vina [45] was used to calculate the binding affinity between the pocket and the ligand, which involves three specific measurements: (i) Vina Score, which provides a direct quantification of binding affinity for generated molecules; (ii) Vina Min, which incorporates a local structure relaxation phase before calculating binding affinity; and (iii) Vina Dock, which employs an additional re-docking procedure to determine the optimal achievable binding affinity. (b) Chemical Property. Chemical properties were measured through five established metrics: (i) the Quantitative Estimation of Drug-likeness (QED) [46], which aggregates multiple desirable molecular properties to assess drug-likeness; (ii) the Synthetic Accessibility (SA) [47], which evaluates the synthetic feasibility of molecular structures; (iii) LogP (octanol-water partition coefficient) is a key physicochemical parameter for drug-likeness evaluation, with optimal values typically falling within the range of -0.4 to 5.6 for orally bioavailable compounds [48]; (iv) the Lipinski score quantifies molecular compliance with Lipinski’s Rule of Five (Ro5) [49], ranging from 0 to 5 based on the number of satisfied criteria; and (v) Diversity is calculated as the average pairwise Tanimoto distance between the molecular fingerprints of generated molecules, measuring structural dissimilarity; (c) Conformation Stability. We evaluated conformation stability from four perspectives: (i) Strain Energy (SE) is used to evaluate the rationality of generated ligand conformation [50], which was calculated by PoseCheck [51]; (ii) the Jensen-Shannon divergence(JSD) of bond length ($\text{JS}_{\text{BL}}$) and bond angle ($\text{JS}_{\text{BA}}$) can reflect the differences in local structures between reference and generated molecules; (iii) Stable Atom Ratio (SAR) is calculated by the fraction of atoms with chemically valid bond configurations, where each atom’s total bond order falls within the permissible range for its element type. The permissible ranges are statistically determined from stable molecules in QM9 [52, 53]. Stable Molecular Ratio (SMR) is calculated by the percentage of generated molecules where all atoms simultaneously satisfy the bond order validity criteria. This represents fully chemically stable molecules. Both metrics employ bond order thresholds fitted to QM9’s bond length distributions, ensuring alignment with known stable molecular configurations. SAR evaluates local atomic stability while SMR assesses global molecular validity [54]; and (iv) Clash Ratio (CR) detects possible clashes in protein-ligand complex [51].
2.4 Model evaluation
Table 1: The comparison of 10,000 generated molecules of MolPIF and baseline models in de novo design scenarios
AR Pk2Mol Target. Decomp. MolCRAFT MolPIF Ref.
Mean Vina Score (↓) -5.75 -5.14 -5.47 -5.19 -6.59 -6.64 -6.36
Median Vina Score (↓) -5.64 -4.7 -6.3 -5.27 -7.04 -7.02 -6.46
Mean Vina Min (↓) -6.18 -6.42 -6.64 -6.03 -7.27 -7.41 -6.71
Median Vina Min (↓) -5.88 -5.82 -6.83 -6 -7.26 -7.28 -6.49
Mean Vina Dock (↓) -6.75 -7.15 -7.8 -7.03 -7.92 -8.09 -7.45
Median Vina Dock (↓) -6.62 -6.79 -7.91 -7.16 -8.01 -8.13 -7.26
Mean QED (↑) 0.51 0.57 0.48 0.51 0.5 0.59 0.48
Mean SA (↑) 0.64 0.76 0.58 0.66 0.69 0.72 0.73
LogP 0.39 1.51 1.36 1.15 1.16 3.26 0.89
Lipinski (↑) 4.75 4.88 4.51 4.49 4.46 4.63 4.27
Diversity (↑) 0.7 0.74 0.72 0.73 0.73 0.72 -
SE 25% (↓) 259 102 369 115 83 65 34
SE 50% (↓) 595 189 1243 421 195 150 107
SE 75% (↓) 2286 374 13871 1424 510 375 196
$\text{JS}_{\text{BL}}$ (↓) 0.4549 0.3721 0.2637 0.2621 0.227 0.2332 -
$\text{JS}_{\text{BA}}$ (↓) 0.5391 0.4275 0.4751 0.4381 0.3686 0.4013 -
SAR (↑) 0.9104 0.8372 0.9492 0.9141 0.9061 0.9643 0.9398
SMR (↑) 0.4649 0.1388 0.42 0.3398 0.3033 0.5366 0.43
CR (↓) 0.2208 0.5585 0.526 0.5098 0.2551 0.2906 0.17
00footnotetext: For each method, we generated approximately 100 molecules per target. Pk2Mol: Pocket2Mol, Target.: TargetDiff, Decomp.: DecompDiff and Ref.: Reference. Vina Score, Vina Min, Vina Dock, SE, $\text{JS}_{\text{BL}}$, $\text{JS}_{\text{BA}}$ and CR are the lower the better (↓); while QED, SA, Lipinski, Diversity, SAR, SMR are the higher the better (↑). The optimal range for LogP is between -0.4 and 5.6. The best two results are highlighted with bold text and underlined text respectively. Note: While the target generation count was set at 100 molecules per method, the actual number of successfully generated molecules may be slightly lower.
2.4.1 Evaluation of common properties for generated molecules
The evaluation results comparing MolPIF with baseline models for de novo design were presented in Table 1. A quantitative assessment was conducted on a dataset comprising 100 molecules per pocket, yielding a total of 10,000 molecules generated by each model. Analysis of the performance metrics revealed that MolPIF excelled in most of the metrics for binding affinity, chemical property and conformation stability, which indicated that MolPIF could effectively learn the binding distribution of protein-ligand complex, and had an excellent ability to generate high-quality molecules.
The experimental results indicated that MolPIF achieved better performance across nearly all evaluation metrics. For binding affinity, when the size of the generated molecules was constrained to be the same as that of the reference molecules—acknowledging that larger molecules tend to achieve better docking score. As demonstrated in Table 1, the molecules generated by MolPIF exhibited superior binding affinity to the target pocket. Notably, MolPIF achieved the lowest mean values across all evaluated metrics: Vina Score (-6.64), Vina Min (-7.41), and Vina Dock (-8.09), outperforming all baseline methods. It was 15.48% better than the best autoregressive baseline, AR and 21.39% better than the diffusion-based baseline, TargetDiff. Even compared to MolCRAFT, which also operates in the parametric space, MolPIF exhibited further enhancements. MolPIF achieved consistently stronger median values than baseline models, with only a negligible difference in Vina Score compared to MolCRAFT (-7.02 vs. -7.04). Importantly, the near-unity ratio (0.82) between the mean Vina Score and Vina Dock further suggested structural consistency between MolPIF’s initial conformations and docked poses. These findings underscored that MolPIF not only generated molecules with optimal binding affinity in their native conformations but also maintained this advantage after rigorous docking refinement. This dual capability implied that the model precisely encoded the pocket’s conformational information while identifying the critical features necessary for forming high-affinity interactions.
MolPIF exhibited superior performance in generating molecules with favorable chemical characteristics. Specifically, it achieved the highest QED (0.59) and the second-highest SA (0.72), trailing only Pocket2Mol. The LogP values of MolPIF-generated molecules also fell within an optimal range (-0.4 to 5.6). In terms of Lipinski’s rule of five compliance, MolPIF ranked third, following AR and Pocket2Mol, while significantly outperforming MolCRAFT-another parametric space-based model. Notably, all evaluated models demonstrated comparable performance in molecular diversity metrics. To the best of our knowledge, these results marked a substantial improvement over existing atom-level molecular generation models reliant on probabilistic path-based methodologies. The findings underscored MolPIF’s enhanced capacity to accurately capture common molecular fragments and essential pharmacophoric features—a task that remains pivotal yet often challenging for atomic-level generative approaches. The robust performance of MolPIF across multiple chemical metrics suggested that its generated molecules exhibited enhanced drug-like properties, potentially offering greater promise for preclinical development.
[…]
The current implementation of MolPIF demonstrated slightly inferior performance in protein-ligand clash ratio compared to AR (by 0.0698) and MolCRAFT (by 0.0355). Notably, when the random masking strategy was omitted during training, MolPIF achieved the second-lowest CR (0.2369). These results implied that simultaneous inclusion of both pocket atoms and partial ligand atoms during training might compromise the model’s ability to accurately learn protein-ligand distances. This phenomenon likely stemmed from the distinct spatial requirements involved: newly incorporated atoms had to simultaneously avoid steric clashes with the binding pocket while satisfying bonding constraints with the supplied ligand atoms. The substantial disparity between these two distance constraints appeared to create a challenging optimization scenario.
2.4.2 Analysis of MolPIF on local geometries
Figure 2: Local geometry analysis. a, Ratio of different-sized rings generated by the model. b, c, d, Jensen-Shannon divergence (JSD) of bond lengths, bond angles and torsion angles distributions between generated molecules and the reference molecules in the test set.
In addition to conventional properties, we evaluated the local geometric characteristics of the generated molecules through (1) statistical comparison of 3- to 8-membered ring frequencies between reference and generated molecules, and (2) substructural level’s Jensen-Shannon divergence (JSD) analysis of key geometric parameters including common covalent bond lengths (CC, C=C, CO, CN, C=N, OP, C=O), bond angles (CCC, C=C=C, CCO, C=C=N, CCN), and torsion angles (CCCC, C=C=C=C, CCOC, CCCO) distributions (Fig.  2). This comprehensive analysis provided insights into the model’s ability to reproduce the precise geometric features of molecular structures.
Fig. 2a presents the occurrence frequencies of 3- to 8-membered rings in both model-generated molecules and reference molecules. MolPIF demonstrated significantly lower proportions of unstable small rings (3- and 4-membered) in its generated molecules. Notably, no 3-membered rings were observed in MolPIF’s outputs, while 4-membered rings occurred at only 0.44% frequency, which was substantially lower than diffusion-model-based baselines. MolPIF exhibited a strong preference for generating 5- and 6-membered rings, which are widely employed in drug design. Specifically, 6-membered rings accounted for 76.88% of generated structures. For larger 7- and 8-membered rings, MolPIF maintained relatively low generation frequencies: 7-membered ring production was comparable to MolCRAFT and slightly higher than AR and Pocket2Mol, while 8-membered rings showed the lowest occurrence among all compared methods (Extended Data Table 3). The AR, constrained by its local topology generation mechanism, tended to overproduce small ring structures. Diffusion-based models TargetDiff and DecompDiff displayed relatively higher frequencies for 4-, 7-, and 8-membered rings but underperformed in generating the most prevalent 6-membered rings, indicating their difficulty in fitting the reference molecular ring distribution during training. In contrast, both MolCRAFT and MolPIF, which operate in parameter space for molecular generation, produced ring distributions that closely approximated the reference molecular profiles.
Fig. 2b-d presents the JSD of bond lengths, bond angles, and torsion angles between molecules generated by MolPIF and baseline methods compared to reference molecules in the test set. Overall, the JSD between covalent bond lengths, bond angles, and torsion angles of model-generated molecules and the reference set followed a consistent trend: autoregressive models >diffusion-based models >parametric generative models. Among parametric generative models, MolPIF demonstrated lower JSD than MolCRAFT for bond lengths of C=C and C=N bonds. The unmasked variant of MolPIF (MolPIF w/o mask) outperformed MolCRAFT in bond length JSD for all bonds except CN and OP. For bond angle distributions, MolPIF showed slightly inferior performance to MolCRAFT, though both MolPIF and its unmasked variant maintained relatively low JSD values overall. In torsion angle distributions, no significant differences were observed except for C=C=C=C—where MolCRAFT performed best—while MolPIF (w/o mask) achieved optimal JSD values for all other torsion angles analyzed. Detailed substructure-level metrics are provided in Extended Data Table 456.
These results indicated that our model effectively learned structural features from reference molecules, generating chemically plausible structures with distributions closely matching the reference data. For the same parametric-space generation framework, MolPIF achieved comparable or superior performance to MolCRAFT in local geometric fidelity.
2.4.3 Case analysis for the performance of MolPIF in de novo molecule generation
For the case study, we selected three representative binding pockets (2V3R, 1L3L, and 6VO5) for de novo molecule generation. Three-dimensional visualization of the generation results is presented in Figure 3. Figure 3a–d illustrate the generation outcomes for pockets 2V3R and 1L3L. The displayed molecules represented the top-performing candidates from 100 generated samples per model, selected according to a comprehensive ranking that considered all evaluation metrics. Quantitative analysis demonstrated that MolPIF-generated molecules exhibited superior performance across multiple metrics under the constraint of a fixed number of atoms. This improved ligand efficiency suggested substantial advantages for downstream optimization and experimental synthesis, substantiating the practical applicability and potential of MolPIF in real-world drug discovery applications.
To evaluate the model’s performance in a biologically relevant context, we applied MolPIF to generate potential inhibitors targeting HAT1. HAT1 is an enzyme that catalyzes the acetylation, which plays a critical role in various physiological processes, and its dysregulation has been closely associated with multiple human diseases, particularly cancers [55]. The generated molecules were then compared with both the reference molecule from the crystal structure and the preliminary screening hits reported in the literature. In this evaluation, MolPIF generated 100 candidate molecules targeting the HAT1 binding site (PDB ID: 6VO5). Remarkably, approximately 50% of these molecules outperformed molecule H9 in key metrics (Figure 3e,f), with the vast majority showing improvement over the reference. Notably, H9 was identified from 100,000 PocketFlow-generated molecules [56] and has been experimentally validated for its biological activity ($IC_{\text{50}}=72.36\mu M$). These results indirectly demonstrated the exceptional efficiency of MolPIF in de novo generation of active molecules as starting points for novel targets.
Figure 3: Case study of generated molecules in de novo generation scenarios. a,b, 3D structures of selected molecules (DecompDiff, MolCRAFT, MolPIF) versus references for targets 2v3r and 1l3l. c,d, Performance comparison of 100 molecules per method against references for 2v3r/1l3l across five metrics: size, Vina score, QED, SA, and SE. e, Vina score/QED/SA distributions for MolPIF-generated molecules (target 6VO5) versus reference and PocketFlow-generated molecule H9. f, 3D structural comparison of reference, H9, and MolPIF molecules for 6VO5. Note: Target generation was 100 molecules per method; actual yields may vary.
2.4.4 Analysis of the prior distribution selection
To conceptually validate the flexibility in prior distribution selection of PIF, we replaced the Gaussian distribution in MolPIF with a Laplace distribution to model the coordinates of atoms. The Laplace distribution, characterized by a sharp peak at the mean and heavier tails compared to the Gaussian distribution, is well-suited for robust statistical modeling [57], signal processing [58], and sparse representations [59]. Following this modification, we retrained the model and evaluated the model on its generative performance using the CrossDocked dataset. For consistency, we generated 100 molecules per binding pocket in the test set for both the Gaussian and Laplace variants of MolPIF and analyzed their performance.
Table 2 presents a comparative analysis of modeling atomic coordinates using Gaussian (MolPIF) and Laplace distributions (MolPIF(La)), along with their respective variants trained without the mask module (MolPIF(w/o mask) and MolPIF(La w/o mask)), evaluated across standard performance metrics. The Gaussian distribution generally demonstrated superior performance across most conventional metrics. Notably, in terms of binding affinity, Laplace-based modeling adversely affected Vina score, Vina min, and Vina dock compared to Gaussian modeling, with mean reductions of 22.14%, 13.23%, and 5.56%, respectively. However, for chemical properties, Laplace-based modeling exhibited slight advantages—MolPIF(La w/o mask) achieved competitive results in QED, SA, LogP, and Lipinski metrics. In molecular diversity, MolPIF(La w/o mask) performed best, attaining a score of 0.75. Regarding molecular conformation stability, MolPIF, MolPIF(w/o mask), and MolPIF(La w/o mask) maintained relatively low SE levels across all quantiles. For bond lengths, MolPIF(La) generated molecules with distributions significantly closer to the reference molecules, yielding a JSD of only 0.1568. In contrast, bond angle distributions showed no notable differences among MolPIF, MolPIF(w/o mask), and MolPIF(La). In SAR, SMR, and CR metrics, Gaussian-based modeling substantially outperformed Laplace-based modeling. However, although the performance was inferior to Gaussian-based MolPIF in conventional metrics, the Laplace-based MolPIF still surpassed numerous baseline models across multiple metrics in Table 1. This observation further demonstrated the robustness of our model when employing different prior distributions.
Regarding the performance of the model in generating molecular substructures, Fig. 4 presents a visualization of the specific distributions for selected representative cases, with detailed data available in Extended Data Tables 78, and 9. The results were compared with MolCRAFT, which demonstrated exceptional performance in molecular substructure generation. For bond lengths, MolPIF showed superior capability in capturing the peak values of reference molecular bond length distributions compared to MolCRAFT. The fitted distributions by MolPIF aligned more closely with the reference bond length distributions (Fig. 4a-c), exhibiting generally lower JSD values (Extended Data Table 7). Notably, the Laplace prior distribution provided a more pronounced advantage, particularly for MolPIF(La), which produced distributions that most closely matched the original among all models. For C=O bonds, MolPIF(La) successfully fit three distinct modes that other models failed to capture (Fig. 4c). In terms of bond angles, MolPIF(La) also demonstrated outstanding performance, achieving the lowest JSD values for C=C=C and CCO distributions compared to the reference molecules, while attaining the second best JSD performance for CCC and CCN (Extended Data Table 8). Taking CCO as an example, MolPIF(La)-generated molecules better approximated the true distribution, with more accurate peak positions and more obvious curve variations than other models (Fig. 4d). Regarding torsion angles, MolPIF(w/o mask) exhibited superior performance, achieving optimal JSD values for CCCC, CCOC, and CCCO, while MolPIF(La) attained the second best performance for C=C=C=C and CCOC (Extended Data Table 9). However, none of the models performed exceptionally well in fitting specific torsion angle modes. For instance, in the case of CCCC, all models only partially captured the distribution modes, with some deviation in peak positions (Fig. 4e).
Overall, when employing a Gaussian distribution as the prior, MolPIF demonstrated superior performance in generating molecules with favorable conventional properties. In contrast, using a Laplace distribution as the prior yielded molecules with substructure distributions that more closely aligned with the reference molecules. We observed an interesting phenomenon: under the Gaussian prior, the mask module enhanced conventional molecular properties but compromised substructure performance. As evident from Fig. 4, MolPIF (w/o mask) exhibited substructure distributions more similar to the reference molecules compared to the standard MolPIF. However, with the Laplace prior, the mask module showed selective improvements - it only benefited certain conventional properties (e.g., Vina Score, Vina Min, $\text{JS}_{\text{BL}}$, and $\text{JS}_{\text{BA}}$) while dramatically enhancing substructure generation. Specifically, the mask module enabled MolPIF(La) to better approximate the reference molecular substructure distributions compared to MolPIF(La w/o mask), demonstrating markedly different effects from its Gaussian counterpart.
We believe these phenomena stem from the critical interplay between the inductive bias of the chosen prior distribution and the specific learning objective introduced by the mask module. One interpretation is that the fundamental difference lies in how Gaussian and Laplace priors model the coordinate space: a Gaussian prior, with its light tails, tends to promote globally smooth and cohesive structures, potentially leading to molecules with conformations that score well on conventional properties like the Vina Score. Conversely, a Laplace prior, characterized by its sharper peak and heavier tails, could be more adept at representing precise, localized geometric features, suggesting an inherent advantage in capturing the rigid arrangements of atoms that define specific substructures—i.e., it aligns better with the local sparsity of substructures (such as the specific geometry of functional groups). When the mask module was introduced, it forced the model to learn a context-aware reconstruction task. In the case of the Gaussian prior, a conflict arose: while learning inter-atomic context improved overall molecular stability and thus conventional properties, the addition of masking noise (randomly obscuring atoms) disrupted the local continuity between atoms. The model failed to effectively reconstruct such partial information, resulting in excessive smoothing of substructure details and consequently a deviation from the reference distribution. In contrast, a synergy occurred with the Laplace prior. The mask module’s objective to ”fill-in-the-blanks” was well complemented by the Laplace prior’s ability to model sharp features. This alignment enabled the model to learn the explicit rules of substructure completion with higher fidelity, leading to more realistic molecules and gains in certain conventional properties. Ultimately, the Gaussian-with-mask setup optimized for global plausibility at the expense of local detail, whereas the Laplace-with-mask setup excelled at local fidelity, proving to be a more effective strategy for generating molecules with substructures that closely match the reference distribution.
These results demonstrated that MolPIF retained strong generative capabilities even with non-conventional priors (e.g., Laplace), achieving competitive performance on conventional benchmarks while exhibiting unique advantages in structural fidelity. The PIF framework’s avoidance of complex closed-form solution derivations—required in diffusion models or BFNs—conferred exceptional flexibility in prior selection. This adaptability opens new possibilities for exploring alternative distributions in SBDD, potentially improving generative performance.
Table 2: The comparison of 10,000 generated molecules of MolPIF variants in de novo design scenarios
MolPIF MolPIF (w/o mask) MolPIF (La) MolPIF (La w/o mask)
Mean Vina Score (↓) -6.64 -6.78 -5.17 -4.75
Median Vina Score (↓) -7.02 -6.99 -6.02 -6.1
Mean Vina Min (↓) -7.41 -7.28 -6.43 -6.28
Median Vina Min (↓) -7.28 -7.18 -6.7 -6.68
Mean Vina Dock (↓) -8.09 -7.9 -7.64 -7.84
Median Vina Dock (↓) -8.13 -7.95 -7.76 -7.9
Mean QED (↑) 0.59 0.55 0.55 0.56
Mean SA (↑) 0.72 0.7 0.7 0.72
LogP 3.26 2.28 2.45 3.42
Lipinski (↑) 4.63 4.49 4.51 4.65
Diversity (↑) 0.72 0.72 0.73 0.75
SE 25% (↓) 65 73 77 81
SE 50% (↓) 150 173 203 183
SE 75% (↓) 375 467 691 417
$\text{JS}_{\text{BL}}$ (↓) 0.2332 0.1965 0.1568 0.2614
$\text{JS}_{\text{BA}}$ (↓) 0.4013 0.3912 0.4024 0.4662
SAR (↑) 0.9643 0.9581 0.9318 0.9365
SMR (↑) 0.5366 0.5338 0.4716 0.5087
CR (↓) 0.2906 0.2369 0.3868 0.3624
00footnotetext: For each method, we generated approximately 100 molecules per target. Note: While the target generation count was set at 100 molecules per method, the actual number of successfully generated molecules may be slightly lower.
Figure 4: Local geometry analysis of reference molecules, MolCRAFT-Generated molecules, and MolPIF Variant-Generated molecules. a-e, distributions of bond lengths (CC, C=C, CO), bond angles (CCO), and torsion angles (CCCC) in molecules generated by the models compared with the test set.
2.4.5 The performance of MolPIF in lead optimization
MolPIF demonstrated applicability not only to de novo molecular generation tasks but also to atom-level lead optimization. This capability enabled arbitrary specification of fixed atoms within given molecules while allowing the model to generate desired substituent regions. Figure 5 presents selected lead optimization cases performed on reference molecules targeting 1umd, 3ZCW, and 6KZZ.
The target 1umd was selected from the CrossDocked test set, with fixed ligand atom selection following the protocol established by CBGBench [54]. Figure 5a displays some molecular structures generated by MolPIF for pocket 1umd during lead optimization. The reference ligand was partitioned into distinct regions for four subtasks: fragment growth, linker design, scaffold hopping, and side-chain decoration. MolPIF exhibited dual functional capabilities: (1) extending existing scaffold groups through Fragment and Side Chain modifications, and (2) integrating discrete fragments into complete molecular structures via Linker and Scaffold operations. Notably, the model accomplished molecular optimization without explicit gradient guidance from property prediction, relying solely on learned structural information. This suggested that the model implicitly captured structural features associated with favorable molecular properties. Such atom-level lead optimization facilitated straightforward modification or replacement of molecular substructures, enabling optimization of physicochemical properties or circumvention of patent restrictions, thereby demonstrating significant practical potential.
We further evaluated MolPIF’s lead optimization performance through two real-world case studies: Kinesin Eg5 [60] (PDB ID: 3ZCW) and E. coli DNA gyrase B [61] (PDB ID: 6KZZ) (Fig. 5b-d). For Kinesin Eg5, we adopted the same side-chain fixation strategy as Delete [62] to maintain consistency. In contrast, for E. coli DNA gyrase B, we relaxed the atomic constraints compared to DeepFrag [63], enabling larger fragment generation to better demonstrate lead optimization potential. Evaluation results confirmed MolPIF’s successful lead optimization in both cases. Subsequent AutoDock Vina scoring revealed that generated molecules exhibited comparable or superior calculated binding affinities relative to original compounds. In scaffold hopping tasks targeting 3ZCW, 25.58% of MolPIF-optimized ligands surpassed the Vina score of the reference ligand(-9.93). Similarly, for fragment growth tasks targeting 6KZZ, 63.33% of MolPIF-optimized ligands surpassed the Vina score of the reference ligand(-8.22).
Across all subtasks involving 1umd, 3ZCW, and 6KZZ, a substantial proportion of the MolPIF-generated molecules demonstrated superior performance compared to the reference molecules in multiple key metrics (Extended Data Tables 10,  11). These findings indicated that MolPIF not only enabled effective structural optimization but also maintained or enhanced molecules’ 3D spatial compatibility and binding capability. Under identical experimental conditions, MolPIF generated a higher proportion of molecules with properties superior to reference compounds compared to Delete and DeepFrag, underscoring its broad applicability and substantial potential in lead optimization tasks.
Figure 5: Case study of generated molecules in lead optimization scenarios. a, 3D structures of selected MolPIF-generated molecules for target 1umd across four optimization scenarios. b,c, 3D structures of MolPIF-generated molecules for targets 3ZCW (scaffold hopping) and 6KZZ (fragment growth), alongside reference molecules. d, Vina score distribution for 100 MolPIF-generated molecules (3ZCW and 6KZZ), compared to reference molecules. Note: Target generation was set at 100 molecules per method, the actual number of successfully generated molecules may be slightly lower.
3 Conclusion
In this study, we propose Parameter Interpolation Flow (PIF), a novel generative framework that operates in the parameter space of probability distributions. PIF constructs a continuous interpolation path between prior and target distributions, which enables smooth transformations while maintaining compatibility with both continuous and discrete data domains. This flexibility, combined with its ability to integrate diverse prior distributions, distinguishes PIF from conventional flow models, diffusion models and bayesian flow networks.
By applying PIF to structure-based drug design, we develop MoIPIF which demonstrates superior performance in generating 3D molecules conditioned on protein binding pockets. Furthermore, we implement a geometry-enhanced training strategy that provides atomic context to assist model training process. Comprehensive evaluations demonstrate the framework’s effectiveness, showing significant improvements in both general molecular properties and local geometries compared to state-of-the-art baselines. Case studies on targets including 2v3r, 1l3l and HAT1 (6VO5) further validate MoIPIF’s ability to generate novel molecules with enhanced binding properties over reference compounds.
The framework’s flexibility is demonstrated through its compatibility with different prior distributions (e.g., Gaussian and Laplace), each offering unique advantages in molecular generation tasks. We further elucidate the mechanistic rationale behind the differential effects induced by integrating Gaussian and Laplace distributions with the mask module. This adaptability, coupled with the model’s robust performance across multiple metrics, establishes MolPIF as a promising new paradigm for generative modeling in computational drug discovery.
For lead optimization scenarios, MoIPIF shows promising capability in modifying specified molecular substructures while preserving critical binding characteristics. In case studies involving 1umd, Kinesin Eg5 (3ZCW) and E. coli DNA gyrase B (6KZZ), the model generates optimized variants, with a substantial proportion of molecules exhibiting enhanced docking scores compared to reference ligands. These results suggest MoIPIF’s potential for integration into structure-based drug discovery pipelines.
Future work may explore extensions to additional distribution types and applications in related molecular design challenges. The success of MoIPIF highlights the potential of parameter-space-based approaches to advance AI-driven drug discovery, enabling efficient exploration of chemical space while preserving structural integrity and chemical validity.
4 Methods
4.1 Definitions and notations
Molecule generation based on receptor structure can be formulated as a conditional generation task. The input is a protein binding site $\mathcal{P}=\{(\textbf{x}_{P}^{(i)},\textbf{v}_{P}^{(i)})\}_{i=1}^{N_{P}}$, which contains $N_{P}$ atoms with each $\textbf{x}_{P}^{(i)}\in\mathbb{R}^{3}$ and $\textbf{v}_{P}^{(i)}\in\mathbb{R}^{D_{P}}$ correspond to atom coordinates and atom features such as element types and amino acid types, respectively. The target output is a binding molecule =$\mathcal{M}=\{(\textbf{x}_{M}^{(i)},\textbf{v}_{M}^{(i)})\}_{i=1}^{N_{M}}$, where $N_{M}$ is the number of atoms in molecule, $\textbf{x}_{M}^{(i)}\in\mathbb{R}^{3}$ and $\textbf{v}_{M}^{(i)}\in\mathbb{R}^{D_{M}}$. For brevity, we denote $\textbf{p}=[\textbf{x}_{P},\textbf{v}_{P}]$ ($\textbf{x}_{P}\in\mathbb{R}^{N_{P}\times 3}$, $\textbf{v}_{P}\in\mathbb{R}^{N_{P}\times D_{P}}$) and $\textbf{m}=[\textbf{x}_{M},\textbf{v}_{M}]$ ($\textbf{x}_{M}\in\mathbb{R}^{N_{M}\times 3}$, $\textbf{v}_{M}\in\mathbb{R}^{N_{M}\times D_{M}}$) as the concatenation of protein binding site and ligand atoms.
4.2 Parameter Interpolation Flow
Parameter Interpolation Flow, the model introduced in this paper, is a flow model that operates in the parameter space. For a given type of probability distribution, its specific form is determined by its parameters. For a set of data points requiring fitting, an appropriate choice of parameters allows the construction of a Dirac distribution:
$p(\textbf{x}\mid\boldsymbol{\theta}(\textbf{x}_{\text{data}}))=\delta(\textbf{x}-\textbf{x}_{\text{data}})\quad$ (1)
Following the idea of flow matching [27], we construct a flow to transform a simple distribution $p(\textbf{x}\mid\boldsymbol{\theta}_{\text{prior}})$ into the desired data distribution $p(\textbf{x}\mid\boldsymbol{\theta}(\textbf{x}_{\text{data}}))$. Unlike conventional flow matching, which constructs the flow in the sample space, we instead build the flow in the parameter space:
$p(\textbf{x}\mid\boldsymbol{\theta}_{t})=p(\textbf{x}\mid f(t)\boldsymbol{\theta}(\textbf{x}_{\text{data}})+(1-f(t))\boldsymbol{\theta}_{\text{prior}})\quad$ (2)
Here, $t\in[0,1]$, and $f(t)$ is a monotonic function satisfying $f(0)=0$ and $f(1)=1$. Thus, $\boldsymbol{\theta}_{t}$ satisfies $\boldsymbol{\theta}_{0}=\boldsymbol{\theta}_{\text{prior}}$ and $\boldsymbol{\theta}_{1}=\boldsymbol{\theta}(\textbf{x}_{\text{data}})$. We draw samples from $p(\textbf{x}\mid\boldsymbol{\theta}_{t})$ and use them as inputs to the model, which is expected to output the parameters $\boldsymbol{\theta}(\textbf{x}_{\text{data}})$ corresponding to the target data distribution.
To evaluate the accuracy of the predicted parameters $\hat{\boldsymbol{\theta}}$, we construct the predicted distribution parameters $\hat{\boldsymbol{\theta}}_{t+\Delta t}$ and the true interpolation distribution parameters $\boldsymbol{\theta}_{t+\Delta t}$ for the next time step $t+\Delta t$, and then compute the KL divergence between them as the loss function:
$\hat{\boldsymbol{\theta}}_{t+\Delta t}=f(t+\Delta t)\hat{\boldsymbol{\theta}}+(1-f(t+\Delta t))\boldsymbol{\theta}_{\text{prior}}\quad$ (3)
$\boldsymbol{\theta}_{t+\Delta t}=f(t+\Delta t)\boldsymbol{\theta}(\textbf{x}_{\text{data}})+(1-f(t+\Delta t))\boldsymbol{\theta}_{\text{prior}}\quad$ (4)
$L_{t}=\mathbb{E}_{p_{\text{data}}}\left[D_{\text{KL}}\left(p(\textbf{x}\mid\boldsymbol{\theta}_{t+\Delta t})\,\|\,p(\textbf{x}\mid\hat{\boldsymbol{\theta}}_{t+\Delta t})\right)\right],\quad t\in[0,1)\quad$ (5)
Given a trained model $\boldsymbol{\Phi}$, the sampling procedure is as follows:
$\hat{\boldsymbol{\theta}}_{t}\rightarrow\hat{\textbf{m}}_{t}\overset{\boldsymbol{\Phi}}{\rightarrow}\hat{\boldsymbol{\theta}}\rightarrow\hat{\boldsymbol{\theta}}_{t+\Delta t}\rightarrow\cdots\quad$ (6)
where $\hat{\boldsymbol{\theta}}_{t}$ represents the model’s prediction of $\boldsymbol{\theta}_{t}$, and $\hat{\textbf{m}}_{t}$ is a sample drawn from the probability distribution parameterized by $\hat{\boldsymbol{\theta}}_{t}$. The sampling process proceeds from $t=0$ to $t=1$, with samples being drawn from the model’s predicted distribution parameterized by $\hat{\boldsymbol{\theta}}_{t}$.
When a specific substructure of the data is fixed as a condition during generation, it can be provided as a conditional input, allowing the model to generate the remaining parts accordingly:
$\hat{\boldsymbol{\theta}}_{\text{cond}}=\boldsymbol{\Phi}(\textbf{m}_{t,\text{cond}}),\quad\textbf{m}_{t,\text{cond}}\sim p(\textbf{m}\mid\boldsymbol{\theta}_{t},\boldsymbol{\theta}_{\text{cond}})\quad$ (7)
The detailed training and sampling algorithms of PIF are presented in Algorithm 1 and Algorithm 2, respectively.
Algorithm 1 Training procedure of PIF
1:probability distribution $p(\textbf{x}|\boldsymbol{\theta})\in\mathcal{P}$, number of steps $n\in\mathbb{N}$, $\gamma\in\mathbb{R}^{+}$, $\boldsymbol{\theta}_{\text{prior}}\in\boldsymbol{\Theta}$, $\textbf{x}_{\text{data}}\in\mathbb{R}^{D}$, neural network $\boldsymbol{\Phi}$, learning rate $\alpha$
2:Sample $i\sim\mathcal{U}\{0,n-1\}$
3:$t\Leftarrow i/n$
4:$\boldsymbol{\theta}(\textbf{x}_{\text{data}})\in\{\textbf{a}\mid p(\textbf{x}|\textbf{a})=\delta(\textbf{x}-\textbf{x}_{\text{data}})\}$
5:$f(t)\Leftarrow 1-\gamma^{t}$
6:$\boldsymbol{\theta}_{t}\Leftarrow f(t)\boldsymbol{\theta}(\textbf{x}_{\text{data}})+(1-f(t))\boldsymbol{\theta}_{\text{prior}}$
7:Sample $\textbf{m}\sim p(\textbf{x}|\boldsymbol{\theta}_{t})$
8:$\hat{\boldsymbol{\theta}}\Leftarrow\boldsymbol{\Phi}(\textbf{m})$
9:$\Delta t\Leftarrow 1/n$
10:$\boldsymbol{\theta}_{t+\Delta t}\Leftarrow f(t+\Delta t)\boldsymbol{\theta}(\textbf{x}_{\text{data}})+(1-f(t+\Delta t))\boldsymbol{\theta}_{\text{prior}}$
11:$\hat{\boldsymbol{\theta}}_{t+\Delta t}\Leftarrow f(t+\Delta t)\hat{\boldsymbol{\theta}}+(1-f(t+\Delta t))\boldsymbol{\theta}_{\text{prior}}$
[…]
13:$\boldsymbol{\Phi}\Leftarrow\boldsymbol{\Phi}-\alpha\nabla_{\boldsymbol{\Phi}}L_{t}$
Algorithm 2 Sampling procedure of PIF
1:probability distribution $p(\textbf{x}|\boldsymbol{\theta})\in\mathcal{P}$, number of steps $n\in\mathbb{N}$, $\gamma\in\mathbb{R}^{+}$, $\boldsymbol{\theta}_{\text{prior}}\in\boldsymbol{\Theta}$, trained neural network $\boldsymbol{\Phi}$
2:$\boldsymbol{\theta}_{0}\Leftarrow\boldsymbol{\theta}_{\text{prior}}$
3:for $i=0$ to $n-1$ do
4:     $t\Leftarrow i/n$
5:     Sample $\textbf{m}\sim p(\textbf{x}|\boldsymbol{\theta}_{t})$
6:     $\hat{\boldsymbol{\theta}}\Leftarrow\boldsymbol{\Phi}(\textbf{m})$
7:     $\Delta t\Leftarrow 1/n$
8:     $f(t)\Leftarrow 1-\gamma^{t}$
9:     $\boldsymbol{\theta}_{t+\Delta t}\Leftarrow f(t+\Delta t)\hat{\boldsymbol{\theta}}+(1-f(t+\Delta t))\boldsymbol{\theta}_{\text{prior}}$
10:end for
11:Sample $\textbf{m}_{1}\sim p(\textbf{x}|\boldsymbol{\theta}_{1})$
12:return $\textbf{m}_{1}$
4.3 Molecule Generation based on PIF
In the context of molecular generation, the application of PIF to the generation of atomic coordinates and types requires specifying their respective distribution types in advance. In this work, we employ Gaussian distribution for atomic coordinates and Dirichlet distribution [64] for atomic types:
$p(\textbf{x})=\mathcal{N}(\textbf{x};\boldsymbol{\mu},\epsilon^{2}\textbf{I})$ (8)
$p(\textbf{v})=\text{Dir}(\textbf{v};\boldsymbol{\alpha})=\frac{1}{B(\boldsymbol{\alpha})}\prod_{i=1}^{K}\textbf{v}_{i}^{\boldsymbol{\alpha}_{i}-1}$ (9)
Here, $\boldsymbol{\mu}$ denotes a three-dimensional vector, $\epsilon^{2}$ is a scalar, and $\boldsymbol{\alpha}$ is a $K$-dimensional vector, where $K$ corresponds to the number of atom types, $B(\boldsymbol{\alpha})$ is the multivariate beta function. Accordingly, the parameters of the two distributions are denoted as $\boldsymbol{\theta}_{\textbf{x}}=(\boldsymbol{\mu},\epsilon^{2})$ and $\boldsymbol{\theta}_{\textbf{v}}=\boldsymbol{\alpha}$, respectively.
To obtain the parameters corresponding to molecular data, we represent them in the form of Dirac distributions associated with the aforementioned two distribution types. In practice, we extend the definition of Gaussian distributions by setting the standard deviation to zero in the distribution parameters of continuous-variable Dirac distributions:
$p(\textbf{x}|\textbf{x}_{\text{data}})=\lim\limits_{\epsilon\to 0^{+}}\mathcal{N}(\textbf{x};\textbf{x}_{\text{data}},\epsilon^{2}\textbf{I})\quad\boldsymbol{\theta}_{\textbf{x},\text{data}}=(\textbf{x}_{\text{data}},0)$ (10)
$p(\textbf{v}|\textbf{v}_{\text{data}})=\text{Dir}(\textbf{v};\textbf{v}_{\text{data}})\quad\boldsymbol{\theta}_{\textbf{v},\text{data}}=\textbf{v}_{\text{data}}=\text{Onehot}(\text{atom type})$ (11)
The formulation of the interpolation process is given as follows:
$\boldsymbol{\theta}_{\textbf{x},t}=f(t)\boldsymbol{\theta}_{\textbf{x},\text{data}}+(1-f(t))\boldsymbol{\theta}_{\textbf{x},\text{prior}}$ (12)
$\boldsymbol{\theta}_{\textbf{v},t}=f(t)\boldsymbol{\theta}_{\textbf{v},\text{data}}+(1-f(t))\boldsymbol{\theta}_{\textbf{v},\text{prior}}$ (13)
$\boldsymbol{\theta}_{\textbf{x},\text{prior}}=(\textbf{0},\epsilon_{0}^{2})$ (14)
$\boldsymbol{\theta}_{\textbf{v},\text{prior}}=(1/K,1/K,\ldots,1/K)$ (15)
$f(t)=1-\gamma^{t}$ (16)
In the above equation, both $\epsilon_{0}$ and $\gamma$ are hyperparameters. The formulation of $f(t)$ encourages the model to focus more on learning fine-grained structures in the molecular data, thereby improving the quality of generation.
Therefore, based on Eq. 5, the loss function can be formulated as follows [65]:
$L_{t-\Delta t,\textbf{x}}=\frac{(1-\gamma^{t})^{2}}{2\gamma^{t}\epsilon_{0}^{2}}\mathbb{E}_{p_{\text{data}}}[\|\hat{\boldsymbol{\theta}}_{\textbf{x}}^{(1)}-\boldsymbol{\theta}_{\textbf{x}}^{(1)}\|^{2}]$ (17)
$L_{t-\Delta t,\textbf{v}}=\mathbb{E}_{p_{\text{data}}}\left[\sum_{i=1}^{K}\ln\frac{\Gamma(\boldsymbol{\theta}_{\textbf{v},t}^{(i)})}{\Gamma(\hat{\boldsymbol{\theta}}_{\textbf{v},t}^{(i)})}+\sum_{i=1}^{K}(\hat{\boldsymbol{\theta}}_{\textbf{v},t}^{(i)}-\boldsymbol{\theta}_{\textbf{v},t}^{(i)})(\psi(\hat{\boldsymbol{\theta}}_{\textbf{v},t}^{(i)})-\psi(1))\right]$ (18)
$L_{t-\Delta t}=\lambda_{\textbf{x}}L_{t-\Delta t,\textbf{x}}+\lambda_{\textbf{v}}L_{t-\Delta t,\textbf{v}}$ (19)
Here, $\Gamma(\textbf{x})$ is the multivariate gamma function, $\psi(\textbf{x})$ is the multivariate digamma function, both $\lambda_{\textbf{x}}$ and $\lambda_{\textbf{v}}$ are hyperparameters to adjust the weight of loss.
To constrain part of the molecular structure during generation, the designated substructure can be incorporated as a conditional input for both coordinates and atom types, enabling the model to generate the rest conditioned on the fixed substructure:
$\boldsymbol{\theta}_{\textbf{x},t,\text{cond}}=\text{Concat}(\boldsymbol{\theta}_{\textbf{x},t},(\textbf{x}_{\text{cond}},0))$ (20)
$\boldsymbol{\theta}_{\textbf{v},t,\text{cond}}=\text{Concat}(\boldsymbol{\theta}_{\textbf{v},t},\textbf{v}_{\text{cond}})$ (21)
4.4 MolPIF based on Laplace distribution
To conceptually validate the distributional flexibility of the PIF model, we modify the coordinate distribution in the MolPIF framework from a Gaussian to a Laplace distribution, and subsequently retrain the model and evaluate its generative performance on the CrossDocked dataset. Notably, similar to using Gaussian distributions as priors for atomic coordinate modeling, we extend the definition of Laplace distributions by setting the parameter $\beta$ to 0 for continuous-variable Dirac distributions. This requires only the following changes in the computational formulation [65]:
$p(\textbf{x})=\text{La}(\textbf{x};\boldsymbol{\alpha},\beta\textbf{I})$ (22)
$p(\textbf{x}|\textbf{x}_{\text{data}})=\lim\limits_{\beta\to 0^{+}}\text{La}(\mathbf{x};\mathbf{x}_{\text{data}},\beta\mathbf{I}),\quad\boldsymbol{\theta}_{\textbf{x},\text{data}}=(\textbf{x}_{\text{data}},0)$ (23)
$\boldsymbol{\theta}_{\textbf{x},\text{prior}}=(\textbf{0},\beta_{0})$ (24)
$L_{t-\Delta t,\textbf{x}}=\mathbb{E}_{p_{\text{data}}}\left[\sum_{i=1}^{3}\left(\exp\left(-\frac{|\hat{\boldsymbol{\theta}}_{\textbf{x}}^{(1,i)}-\boldsymbol{\theta}_{\textbf{x}}^{(1,i)}|}{\beta_{0}\gamma^{t}}\right)+\frac{|\hat{\boldsymbol{\theta}}_{\textbf{x}}^{(1,i)}-\boldsymbol{\theta}_{\textbf{x}}^{(1,i)}|}{\beta_{0}\gamma^{t}}\right)\right]$ (25)
Similarly, both $\beta_{0}$ and $\gamma$ are hyperparameters.
4.5 Implementation Details
The chemical distribution analysis of molecules within binding pockets was carried out using the CrossDocked dataset. The dataset underwent rigorous preprocessing and splitting procedures, as outlined in prior studies [66, 67], retaining only diverse, high-quality docking poses. This process resulted in 99,900 training pairs and 100 validation pairs. In this study, we set $\epsilon_{0}=1$, $\gamma=0.009$, $\beta_{0}=1$, $P_{m}=0.3$ and $P_{am}=0.3$. To balance efficiency and precision, the number of sample steps was fixed at 100. For lead optimization tasks, the prior parameters were initialized based on the coordinates and atom types of the unfixed regions. The model was trained on a single NVIDIA 4090 GPU for  24 hours.
For the model architecture of MolPIF, the block was built upon the UniTransformer, as utilized in TargetDiff, which ensured equivariance and enabled effective encoding of spatial features for both protein binding sites and ligand molecules. In the graph representations of proteins and ligands, atoms were represented as nodes, with edges determined using the K-Nearest Neighbor (KNN) method [68].
5 Data availability
The evaluation dataset CrossDocked2020 [36] was obtained from the prior study TargetDiff [42] and is available at https://drive.google.com/drive/folders/1j21cc7-97TedKh_El5E34yI8o5ckI7eK. The molecular files used for model testing were sourced from MolCRAFT [29] and can be downloaded at https://drive.google.com/drive/folders/1A3Mthm9ksbfUnMCe5T2noGsiEV1RfChH. Our model weights, configuration files, and generated molecules are publicly available at https://drive.google.com/drive/folders/1VBGnHyThNHpdaLgppOeKCKomwfL6oXde.
6 Code availability
The code of MolPIF is freely available at https://github.com/BLEACH366/MolPIF.
Acknowledgements
This work was supported by Shanghai Rising-Star Program (23QD1400600) and National Key Research and Development Program of China (2022YFC3400504).
Appendix A Toy Data Results
To evaluate our model’s generalization capability, we conducted experiments on several 2D synthetic datasets. These included: (1) the swissroll and swissroll+moons datasets representing continuous distributions, and (2) sparse and dense chessboard datasets simulating discrete distributions. We replaced SLDM’s moons dataset with the more complex swissroll+moons variant and introduced a dense chessboard configuration to further increase distribution complexity, challenging the model’s generalization capability. All datasets contained 100,000 samples. Except for the dense chessboard experiments, hyperparameters matched those in SLDM, following https://github.com/albarji/toy-diffusion/; for dense chessboard, we increased training epochs to 10,000 and diffusion steps to 500 to ensure convergence (compared to 100 epochs/40 steps for swissroll and swissroll+moons, and 600 epochs/100 steps for sparse chessboard). All experiments used a 6-layer MLP with a batch size of 2,048, optimized using Adam (lr = 0.001). For SLDM, we disabled temperature control during sampling, consistent with its original implementation. As shown in Fig. 6, our model generated samples that aligned well with the original distributions, exhibiting fewer outliers and reasonable coverage compared to baseline methods. These results suggested our approach had the ability to effectively handle both continuous and discrete patterns.
Figure 6: Generative performance comparison on toy datasets. a-d, Performance of the DDPM on the swissroll, swissroll+moons, sparse chessboard, and dense chessboard datasets. e-h, Performance of the flow matching on the four datasets. i-l, Performance of the SLDM on the four datasets. m-p, Performance of the BFN on the four datasets. q-t, Performance of the PIF on the four datasets.
Appendix B Extended Data
Table 3: Ratio of different-sized rings generated by models
3 4 5 6 7 8
AR 0.3086 0.0030 0.1556 0.4926 0.0190 0.0082
Pocket2Mol 0.0012 0.0002 0.1626 0.7983 0.0259 0.0034
TargetDiff 0.0000 0.0270 0.2971 0.4896 0.1170 0.0259
DecompDiff 0.0264 0.0391 0.3425 0.4396 0.1135 0.0178
MolCRAFT 0.0000 0.0022 0.2310 0.6986 0.0540 0.0062
MolPIF 0.0000 0.0044 0.1597 0.7688 0.0565 0.0034
MolPIF(w/o mask) 0.0000 0.0045 0.1918 0.7269 0.0614 0.0059
Test set 0.0172 0.0000 0.2961 0.6609 0.0086 0.0000
Table 4: JSD of bond lengths between reference and generated molecules
CC C=C CO CN C=N OP C=O
AR 0.6098 0.4503 0.4904 0.4722 0.5514 0.6762 0.5563
Pocket2Mol 0.4937 0.4144 0.4524 0.4222 0.4840 0.5257 0.5095
TargetDiff 0.3690 0.2634 0.4208 0.3623 0.2349 0.4430 0.4608
DecompDiff 0.3265 0.2378 0.3467 0.3374 0.2344 0.4524 0.3903
MolCRAFT 0.3206 0.3231 0.3536 0.3010 0.2497 0.3410 0.3339
MolPIF 0.3865 0.1978 0.4104 0.3425 0.2068 0.3659 0.3351
MolPIF(w/o mask) 0.2885 0.2257 0.3475 0.3161 0.2216 0.3505 0.3278
Table 5: JSD of bond angles between reference and generated molecules
CCC C=C=C CCO C=C=N CCN
AR 0.3721 0.5719 0.4770 0.5366 0.4474
Pocket2Mol 0.3799 0.4804 0.4753 0.5060 0.4432
TargetDiff 0.3450 0.2828 0.4396 0.4537 0.4374
DecompDiff 0.3383 0.2549 0.3898 0.4288 0.4179
MolCRAFT 0.3015 0.1741 0.3473 0.4508 0.3796
MolPIF 0.3716 0.2078 0.4349 0.4323 0.4178
MolPIF(w/o mask) 0.3073 0.2219 0.3751 0.4509 0.4099
Table 6: JSD of torsion angles between reference and generated molecules
CCCC C=C=C=C CCOC CCCO
AR 0.3778 0.7130 0.3736 0.4270
Pocket2Mol 0.3220 0.5339 0.3683 0.4009
TargetDiff 0.3134 0.3730 0.3416 0.4030
DecompDiff 0.3455 0.2403 0.3423 0.4113
MolCRAFT 0.2818 0.1555 0.3417 0.3977
MolPIF 0.2901 0.3022 0.3417 0.3911
MolPIF(w/o mask) 0.2676 0.2667 0.3337 0.3853
Table 7: JSD of bond lengths between reference and the molecules generated by MolCRAFT and MolPIF variants
CC C=C CO CN C=N OP C=O
MolCRAFT 0.3206 0.3231 0.3536 0.3010 0.2497 0.3410 0.3339
MolPIF 0.3865 0.1978 0.4104 0.3425 0.2068 0.3659 0.3351
MolPIF(w/o mask) 0.2885 0.2257 0.3475 0.3161 0.2216 0.3505 0.3278
MolPIF(La) 0.2642 0.1773 0.2527 0.2796 0.2269 0.3213 0.2889
MolPIF(La w/o mask) 0.4322 0.1555 0.4714 0.3421 0.2137 0.4398 0.3255
Table 8: JSD of bond angles between reference and the molecules generated by MolCRAFT and MolPIF variants
CCC C=C=C CCO C=C=N CCN
MolCRAFT 0.3015 0.1741 0.3473 0.4508 0.3796
MolPIF 0.3716 0.2078 0.4349 0.4323 0.4178
MolPIF(w/o mask) 0.3073 0.2219 0.3751 0.4509 0.4099
MolPIF(La) 0.3025 0.1664 0.3290 0.4670 0.4058
MolPIF(La w/o mask) 0.4747 0.2675 0.5551 0.4785 0.4520
Table 9: JSD of torsion angles between reference and the molecules generated by MolCRAFT and MolPIF variants
CCCC C=C=C=C CCOC CCCO
MolCRAFT 0.2818 0.1555 0.3417 0.3977
MolPIF 0.2901 0.3022 0.3417 0.3911
MolPIF(w/o mask) 0.2676 0.2667 0.3337 0.3853
MolPIF(La) 0.2953 0.1734 0.3388 0.4019
MolPIF(La w/o mask) 0.3600 0.2559 0.3555 0.3938
Table 10: The comparison of 100 generated molecules of MolPIF in lead optimization scenarios
1umd 3ZCW 6KZZ
Metric Frag Linker Scaffold Side chain Ref. Scaffold Ref Frag Ref.
Atoms num 26 26 26 26 26 33 33 23.27 25
Mean Vina Score (↓) -8.42 -9.40 -8.58 -8.37 -8.88 -9.29 -9.93 -8.36 -8.22
Mean Vina Min (↓) -8.64 -9.46 -8.95 -8.75 -8.84 -9.84 -10.10 -8.60 -8.78
Mean Vina Dock (↓) -9.02 -9.67 -9.33 -9.33 -9.39 -10.37 -10.24 -8.72 -9.12
Mean QED (↑) 0.32 0.31 0.28 0.48 0.44 0.48 0.41 0.65 0.58
Mean SA (↑) 0.58 0.57 0.58 0.56 0.66 0.67 0.80 0.81 0.88
LogP 1.08 0.76 0.62 0.35 1.72 5.32 5.11 1.80 2.52
Lipinski (↑) 4.51 4.48 4.28 4.67 5.00 4.33 4.00 4.98 5.00
SE 25% (↓) 350.39 334.74 401.24 315.30 - 252.25 - 144.90 -
SE 50% (↓) 470.68 464.40 639.84 868.76 - 332.79 - 224.68 -
SE 75% (↓) 1061.24 2152.76 2739.69 57074.30 - 563.43 - 280.16 -
SE - - - - 276.19 - 55.86 - 224.81
CR (↓) 17.54 15.06 19.60 19.49 13.00 3.78 0.00 3.85 2.00
00footnotetext: Frag: Fragment growth, Linker: Linker design, Scaffold: Scaffold hopping, Side chain: Side-chain decoration and Ref.: Reference.
Table 11: The proportion of MolPIF-generated molecules outperforming reference compounds in key metrics during lead optimization
1umd 3ZCW 6KZZ
Metric Frag Linker Scaffold Side chain Scaffold Frag
Vina Score 0.33 0.76 0.44 0.42 0.26 0.63
Vina Min 0.42 0.84 0.57 0.54 0.40 0.43
Vina Dock 0.28 0.69 0.43 0.51 0.59 0.23
QED 0.06 0.03 0.04 0.53 0.71 0.77
SA 0.01 0.02 0.03 0.05 0.07 0.22
Lipinski 0.58 0.63 0.46 0.69 1.00 0.98
SE 0.03 0.03 0.06 0.20 0.00 0.50
CR 0.11 0.15 0.00 0.13 0.10 0.28
00footnotetext: Frag: Fragment growth, Linker: Linker design, Scaffold: Scaffold hopping and Side chain: Side-chain decoration.

### v2

@@
 By contrast, such flexibility serves as the critical advantage that sets flow-matching models apart from other paradigms.
-Therefore, to overcome the respective limitations of flow matching and BFNs, we integrate their core concepts and introduce a novel and versatile parameter interpolation flow (PIF) model, as illustrated in Fig.
+Therefore, to overcome the respective limitations of flow matching and BFNs, we introduce a novel and versatile parameter interpolation flow (PIF) model, as illustrated in Fig.
 1.
@@
 2.4.2 Analysis of MolPIF on local geometries
-Figure 2: Local geometry analysis.
-a, Ratio of different-sized rings generated by the model.
-b, c, d, Jensen-Shannon divergence (JSD) of bond lengths, bond angles and torsion angles distributions between generated molecules and the reference molecules in the test set.
-In addition to conventional properties, we evaluated the local geometric characteristics of the generated molecules through (1) statistical comparison of 3- to 8-membered ring frequencies between reference and generated molecules, and (2) substructural level’s Jensen-Shannon divergence (JSD) analysis of key geometric parameters including common covalent bond lengths (CC, C=C, CO, CN, C=N, OP, C=O), bond angles (CCC, C=C=C, CCO, C=C=N, CCN), and torsion angles (CCCC, C=C=C=C, CCOC, CCCO) distributions (Fig.
-2).
+In addition to conventional properties, we evaluated the local geometric characteristics of the generated molecules through (1) statistical comparison of 3- to 8-membered ring frequencies between reference and generated molecules (Table 2) , and (2) substructural level’s Jensen-Shannon divergence (JSD) analysis of key geometric parameters including common covalent bond lengths (CC, C=C, CO, CN, C=N, OP, C=O), bond angles (CCC, C=C=C, CCO, C=C=N, CCN), and torsion angles (CCCC, C=C=C=C, CCOC, CCCO) distributions.
 This comprehensive analysis provided insights into the model’s ability to reproduce the precise geometric features of molecular structures.
-Fig.
-2a presents the occurrence frequencies of 3- to 8-membered rings in both model-generated molecules and reference molecules.
+Table 2 presents the occurrence frequencies of 3- to 8-membered rings in both model-generated molecules and reference molecules.
 MolPIF demonstrated significantly lower proportions of unstable small rings (3- and 4-membered) in its generated molecules.
@@
 Specifically, 6-membered rings accounted for 76.88% of generated structures.
-For larger 7- and 8-membered rings, MolPIF maintained relatively low generation frequencies: 7-membered ring production was comparable to MolCRAFT and slightly higher than AR and Pocket2Mol, while 8-membered rings showed the lowest occurrence among all compared methods (Extended Data Table 3).
+For larger 7- and 8-membered rings, MolPIF maintained relatively low generation frequencies: 7-membered ring production was comparable to MolCRAFT and slightly higher than AR and Pocket2Mol, while 8-membered rings showed the lowest occurrence among all compared methods.
 The AR, constrained by its local topology generation mechanism, tended to overproduce small ring structures.
@@
 In contrast, both MolCRAFT and MolPIF, which operate in parameter space for molecular generation, produced ring distributions that closely approximated the reference molecular profiles.
-Fig.
-2b-d presents the JSD of bond lengths, bond angles, and torsion angles between molecules generated by MolPIF and baseline methods compared to reference molecules in the test set.
-Overall, the JSD between covalent bond lengths, bond angles, and torsion angles of model-generated molecules and the reference set followed a consistent trend: autoregressive models >diffusion-based models >parametric generative models.
-Among parametric generative models, MolPIF demonstrated lower JSD than MolCRAFT for bond lengths of C=C and C=N bonds.
-The unmasked variant of MolPIF (MolPIF w/o mask) outperformed MolCRAFT in bond length JSD for all bonds except CN and OP.
-For bond angle distributions, MolPIF showed slightly inferior performance to MolCRAFT, though both MolPIF and its unmasked variant maintained relatively low JSD values overall.
-In torsion angle distributions, no significant differences were observed except for C=C=C=C—where MolCRAFT performed best—while MolPIF (w/o mask) achieved optimal JSD values for all other torsion angles analyzed.
-Detailed substructure-level metrics are provided in Extended Data Table 456.
+For the JSD of bond lengths, bond angles, and torsion angles between molecules generated by MolPIF and baseline methods compared to reference molecules in the test set.
+Overall, the JSD between covalent bond lengths, bond angles, and torsion angles of model-generated molecules and the reference set followed a consistent trend: autoregressive models >diffusion-based models >parametric generative models, which aligned with the conclusions demonstrated in MolCRAFT [29].
+The specific performance of the superior-performing parametric generative models regarding the JSD of bond lengths, bond angles, and torsion angles will be discussed in detail in Section 2.4.4.
 These results indicated that our model effectively learned structural features from reference molecules, generating chemically plausible structures with distributions closely matching the reference data.
 For the same parametric-space generation framework, MolPIF achieved comparable or superior performance to MolCRAFT in local geometric fidelity.
+Table 2: Ratio of different-sized rings generated by models
+3 4 5 6 7 8
+AR 0.3086 0.0030 0.1556 0.4926 0.0190 0.0082
+Pocket2Mol 0.0012 0.0002 0.1626 0.7983 0.0259 0.0034
+TargetDiff 0.0000 0.0270 0.2971 0.4896 0.1170 0.0259
+DecompDiff 0.0264 0.0391 0.3425 0.4396 0.1135 0.0178
+MolCRAFT 0.0000 0.0022 0.2310 0.6986 0.0540 0.0062
+MolPIF 0.0000 0.0044 0.1597 0.7688 0.0565 0.0034
+MolPIF(w/o mask) 0.0000 0.0045 0.1918 0.7269 0.0614 0.0059
+Test set 0.0172 0.0000 0.2961 0.6609 0.0086 0.0000
 2.4.3 Case analysis for the performance of MolPIF in de novo molecule generation
@@
 Regarding the performance of the model in generating molecular substructures, Fig.
-4 presents a visualization of the specific distributions for selected representative cases, with detailed data available in Extended Data Tables 78, and 9.
+3 presents a visualization of the specific distributions for selected representative cases, with detailed data available in Extended Data Tables 45, and 6.
 The results were compared with MolCRAFT, which demonstrated exceptional performance in molecular substructure generation.
@@
 The fitted distributions by MolPIF aligned more closely with the reference bond length distributions (Fig.
-4a-c), exhibiting generally lower JSD values (Extended Data Table 7).
+3a-c), exhibiting generally lower JSD values (Extended Data Table 4).
 Notably, the Laplace prior distribution provided a more pronounced advantage, particularly for MolPIF(La), which produced distributions that most closely matched the original among all models.
 For C=O bonds, MolPIF(La) successfully fit three distinct modes that other models failed to capture (Fig.
-4c).
+3c).
 In terms of bond angles, MolPIF(La) also demonstrated outstanding performance, achieving the lowest JSD values for C=C=C and CCO distributions compared to the reference molecules, while attaining the second best JSD performance for CCC and CCN (Extended Data Table 5).
 Taking CCO as an example, MolPIF(La)-generated molecules better approximated the true distribution, with more accurate peak positions and more obvious curve variations than other models (Fig.
-4d).
+3d).
 Regarding torsion angles, MolPIF(w/o mask) exhibited superior performance, achieving optimal JSD values for CCCC, CCOC, and CCCO, while MolPIF(La) attained the second best performance for C=C=C=C and CCOC (Extended Data Table 6).
@@
 For instance, in the case of CCCC, all models only partially captured the distribution modes, with some deviation in peak positions (Fig.
-4e).
+3e).
 Overall, when employing a Gaussian distribution as the prior, MolPIF demonstrated superior performance in generating molecules with favorable conventional properties.
@@
 As evident from Fig.
-4, MolPIF (w/o mask) exhibited substructure distributions more similar to the reference molecules compared to the standard MolPIF.
+3, MolPIF (w/o mask) exhibited substructure distributions more similar to the reference molecules compared to the standard MolPIF.
 However, with the Laplace prior, the mask module showed selective improvements - it only benefited certain conventional properties (e.g., Vina Score, Vina Min, $\text{JS}_{\text{BL}}$, and $\text{JS}_{\text{BA}}$) while dramatically enhancing substructure generation.
@@
 coli DNA gyrase B [61] (PDB ID: 6KZZ) (Fig.
-5b-d).
+4b-d).
 For Kinesin Eg5, we adopted the same side-chain fixation strategy as Delete [62] to maintain consistency.
@@
 Similarly, for fragment growth tasks targeting 6KZZ, 63.33% of MolPIF-optimized ligands surpassed the Vina score of the reference ligand(-8.22).
-Across all subtasks involving 1umd, 3ZCW, and 6KZZ, a substantial proportion of the MolPIF-generated molecules demonstrated superior performance compared to the reference molecules in multiple key metrics (Extended Data Tables 10,  11).
+Across all subtasks involving 1umd, 3ZCW, and 6KZZ, a substantial proportion of the MolPIF-generated molecules demonstrated superior performance compared to the reference molecules in multiple key metrics (Extended Data Tables 7,  8).
 These findings indicated that MolPIF not only enabled effective structural optimization but also maintained or enhanced molecules’ 3D spatial compatibility and binding capability.
@@
 As shown in Fig.
-6, our model generated samples that aligned well with the original distributions, exhibiting fewer outliers and reasonable coverage compared to baseline methods.
+5, our model generated samples that aligned well with the original distributions, exhibiting fewer outliers and reasonable coverage compared to baseline methods.
 These results suggested our approach had the ability to effectively handle both continuous and discrete patterns.
@@
 Appendix B Extended Data
-Table 3: Ratio of different-sized rings generated by models
-3 4 5 6 7 8
-AR 0.3086 0.0030 0.1556 0.4926 0.0190 0.0082
-Pocket2Mol 0.0012 0.0002 0.1626 0.7983 0.0259 0.0034
-TargetDiff 0.0000 0.0270 0.2971 0.4896 0.1170 0.0259
-DecompDiff 0.0264 0.0391 0.3425 0.4396 0.1135 0.0178
-MolCRAFT 0.0000 0.0022 0.2310 0.6986 0.0540 0.0062
-MolPIF 0.0000 0.0044 0.1597 0.7688 0.0565 0.0034
-MolPIF(w/o mask) 0.0000 0.0045 0.1918 0.7269 0.0614 0.0059
-Test set 0.0172 0.0000 0.2961 0.6609 0.0086 0.0000
-Table 4: JSD of bond lengths between reference and generated molecules
-CC C=C CO CN C=N OP C=O
-AR 0.6098 0.4503 0.4904 0.4722 0.5514 0.6762 0.5563
-Pocket2Mol 0.4937 0.4144 0.4524 0.4222 0.4840 0.5257 0.5095
-TargetDiff 0.3690 0.2634 0.4208 0.3623 0.2349 0.4430 0.4608
-DecompDiff 0.3265 0.2378 0.3467 0.3374 0.2344 0.4524 0.3903
-MolCRAFT 0.3206 0.3231 0.3536 0.3010 0.2497 0.3410 0.3339
-MolPIF 0.3865 0.1978 0.4104 0.3425 0.2068 0.3659 0.3351
-MolPIF(w/o mask) 0.2885 0.2257 0.3475 0.3161 0.2216 0.3505 0.3278
-Table 5: JSD of bond angles between reference and generated molecules
-CCC C=C=C CCO C=C=N CCN
-AR 0.3721 0.5719 0.4770 0.5366 0.4474
-Pocket2Mol 0.3799 0.4804 0.4753 0.5060 0.4432
-TargetDiff 0.3450 0.2828 0.4396 0.4537 0.4374
-DecompDiff 0.3383 0.2549 0.3898 0.4288 0.4179
-MolCRAFT 0.3015 0.1741 0.3473 0.4508 0.3796
-MolPIF 0.3716 0.2078 0.4349 0.4323 0.4178
-MolPIF(w/o mask) 0.3073 0.2219 0.3751 0.4509 0.4099
-Table 6: JSD of torsion angles between reference and generated molecules
-CCCC C=C=C=C CCOC CCCO
-AR 0.3778 0.7130 0.3736 0.4270
-Pocket2Mol 0.3220 0.5339 0.3683 0.4009
-TargetDiff 0.3134 0.3730 0.3416 0.4030
-DecompDiff 0.3455 0.2403 0.3423 0.4113
-MolCRAFT 0.2818 0.1555 0.3417 0.3977
-MolPIF 0.2901 0.3022 0.3417 0.3911
-MolPIF(w/o mask) 0.2676 0.2667 0.3337 0.3853
 Table 4: JSD of bond lengths between reference and the molecules generated by MolCRAFT and MolPIF variants

### v3

@@
 As a result, MolPIF achieves superior performance in overall quality of the generated molecules.
-Empirical evaluations conducted on the CrossDocked2020 dataset [36] demonstrate that MolPIF has comprehensive advantages across four key dimensions: (1) Superior de novo generation capability, producing molecules with superior binding properties and chemical validity; (2) Accurate geometric reproduction of molecular structural distributions, including rings, bond lengths, and bond angles; (3) Flexible adaptation to different prior distributions, where we systematically compared and analyzed the advantages and limitations of using Gaussian versus Laplace distributions as priors for modeling atomic coordinates; (4) Effective lead optimization, demonstrating robust performance in enhancing drug candidate properties.
+Empirical evaluations conducted on the CrossDocked2020 dataset [36] demonstrate that MolPIF has comprehensive advantages across five key dimensions: (1) Superior de novo generation capability, producing molecules with superior binding properties and chemical validity; (2) Accurate geometric reproduction of molecular structural distributions, including rings, bond lengths, and bond angles; (3) Comprehensive chemical space modeling, with substantial coverage of 2D structural features and accurate 3D conformational distribution, while extending shape diversity beyond the reference distribution; (4) Flexible adaptation to different prior distributions, where we systematically compared and analyzed the advantages and limitations of using Gaussian versus Laplace distributions as priors for modeling atomic coordinates; (5) Effective lead optimization, demonstrating robust performance in enhancing drug candidate properties.
 Our main contributions can be summarized as follows:
@@
 Test set 0.0172 0.0000 0.2961 0.6609 0.0086 0.0000
+2.4.3 Analysis of MolPIF on chemical space distribution
+In contrast to analyzing the local geometry of molecules generated by MolPIF, the examination of their chemical space distribution provided a more macroscopic perspective.
+Inspired by PMDM [55], we employed both 2D and 3D molecular descriptors—the Extended-Connectivity Fingerprints(ECFP), RDKit, and USRCAT ((Ultrafast Shape Recognition with CREDO Atom Types) [56]—to represent the chemical space of generated molecules and reference test set molecules, with particular emphasis on structural topology.
+Specifically, we adopted the ECFP, which implement the Morgan [57] algorithm to assign unique atom identifiers.
+These fingerprints encode atomic environments by integrating atom types (e.g., connectivity), chemical features (e.g., hydrogen bond donors/acceptors), and neighboring atomic contexts.
+The RDKit fingerprint, conceptually derived from the Daylight fingerprint, quantified 2D molecular substructures by evaluating atom and bond types.
+Conversely, USRCAT enhanced the Ultrafast Shape Recognition (USR) algorithm by integrating pharmacophoric descriptors to characterize 3D molecular shape.
+The chemical space distribution visualized by t-SNE [58] was presented in Fig.
+2a–c.
+While MolPIF-generated molecules basically covered the test set’s chemical space in 2D substructure representation (Fig.
+2a,b), they showed superior density alignment in 3D conformational space (Fig.
+2c).
+The generated molecules’ high-density regions precisely matched the test set’s concentrated areas in 3D space, demonstrating MolPIF’s ability to capture authentic conformational distributions.
+This accurate density reproduction in 3D space was particularly notable given molecular conformation complexity.
+The visualization confirmed that MolPIF maintained comprehensive 2D feature coverage while achieving physically meaningful 3D distributions that highlight the most relevant conformational regions observed experimentally.
+Figure 2: Visualization of chemical space distribution and molecular shape characteristics.
+a-c, Chemical space distributions of molecules visualized via t-SNE in two-dimensional space, based on Morgan (a), RDKit (b), and USRCAT (c) fingerprints, respectively.
+d,e, Shape distributions are compared between generated molecules (d) and reference set molecules (e), represented using Normalized Principal Moment of Inertia ratios (NPR).
+f, Statistical comparison of Plane of Best Fit (PBF) descriptor values between generated molecules (n=10,000) and reference set (n=100), with box plots showing median (center line), interquartile range (box limits), 1.5×IQR whiskers, and extreme values.
+To further analyze the 3D shape distribution of generated molecules, we employed molecular descriptors to characterize molecular structures beyond conventional fingerprints.
+Following the methodology used in PMDM, we utilized two widely recognized molecular descriptors: Principal Moments of Inertia (PMI) [59] and Plane of Best Fit (PBF) [60], which enable comprehensive investigation of molecular shapes.
+The PMI analysis quantitatively characterized molecular geometries by classifying them into rod-shaped, disc-shaped, or sphere-shaped configurations based on their inertial properties.
+Meanwhile, the PBF approach determined the optimal fitting plane through all heavy atoms in a given molecular conformation and subsequently computed the average deviation of these atoms from the reference plane, providing a complementary measure of molecular planarity.
+Fig.
+2d,e visualize the Normalized Principal Moment of Inertia ratios (NPR) on a ternary plot, where proximity to each corner reflects the dominance of rod-, disc-, or sphere-like morphologies.
+Notably, MolPIF-generated molecules exhibited a distribution closely aligned with the test set, with both populations predominantly clustered near the rod-like vertex.
+Importantly, MolPIF further extended the shape diversity beyond the reference’s coverage, as evidenced by generated samples reaching the disc- and sphere-like regions—areas sparsely represented (or absent) in the original test distribution.
+This demonstrated MolPIF’s dual capability: (1) faithfully learning the 3D molecular shape distribution of the datasets, and (2) exploring novel molecular structures.
+Moreover, as illustrated in Fig.
+2f, the PBF value distribution of molecules generated by MolPIF showed agreement with that of the test set, demonstrating the model’s capability to accurately learn and reproduce the spatial planarity characteristics of authentic molecular conformations.
+This precise modeling of PBF metrics indicated that the model had successfully captured the intrinsic relationship between the spatial arrangement of heavy atoms and molecular planarity.
+MolPIF effectively captured and extended the chemical space of reference molecules, achieving strong coverage in both 2D substructure and 3D conformational representations.
+While maintaining fidelity to dominant structural trends, it also explored novel shape diversity beyond the original distribution, as demonstrated by PMI and PBF analysis.
+This balance between accurate learning and innovative generation highlighted its potential for drug discovery applications requiring precise 3D molecular design.
 2.4.4 Case analysis for the performance of MolPIF in de novo molecule generation
 For the case study, we selected three representative binding pockets (2V3R, 1L3L, and 6VO5) for de novo molecule generation.
-Three-dimensional visualization of the generation results is presented in Figure 2.
-Figure 2a–d illustrate the generation outcomes for pockets 2V3R and 1L3L.
+Three-dimensional visualization of the generation results is presented in Fig.
+3.
+Fig.
+3a–d illustrate the generation outcomes for pockets 2V3R and 1L3L.
 The displayed molecules represented the top-performing candidates from 100 generated samples per model, selected according to a comprehensive ranking that considered all evaluation metrics.
@@
 In this evaluation, MolPIF generated 100 candidate molecules targeting the HAT1 binding site (PDB ID: 6VO5).
-Remarkably, approximately 50% of these molecules outperformed molecule H9 in key metrics (Figure 2e,f), with the vast majority showing improvement over the reference.
+Remarkably, approximately 50% of these molecules outperformed molecule H9 in key metrics (Fig.
+3e,f), with the vast majority showing improvement over the reference.
 Notably, H9 was identified from 100,000 PocketFlow-generated molecules [62] and has been experimentally validated for its biological activity ($IC_{\text{50}}=72.36\mu M$).
@@
 The fitted distributions by MolPIF aligned more closely with the reference bond length distributions (Fig.
-3a-c), exhibiting generally lower JSD values (Extended Data Table 4).
+4a-c), exhibiting generally lower JSD values (Extended Data Table 4).
 Notably, the Laplace prior distribution provided a more pronounced advantage, particularly for MolPIF(La), which produced distributions that most closely matched the original among all models.
 For C=O bonds, MolPIF(La) successfully fit three distinct modes that other models failed to capture (Fig.
-3c).
+4c).
 In terms of bond angles, MolPIF(La) also demonstrated outstanding performance, achieving the lowest JSD values for C=C=C and CCO distributions compared to the reference molecules, while attaining the second best JSD performance for CCC and CCN (Extended Data Table 5).
 Taking CCO as an example, MolPIF(La)-generated molecules better approximated the true distribution, with more accurate peak positions and more obvious curve variations than other models (Fig.
-3d).
+4d).
 Regarding torsion angles, MolPIF(w/o mask) exhibited superior performance, achieving optimal JSD values for CCCC, CCOC, and CCCO, while MolPIF(La) attained the second best performance for C=C=C=C and CCOC (Extended Data Table 6).
@@
 For instance, in the case of CCCC, all models only partially captured the distribution modes, with some deviation in peak positions (Fig.
-3e).
+4e).
 Overall, when employing a Gaussian distribution as the prior, MolPIF demonstrated superior performance in generating molecules with favorable conventional properties.
@@
 As evident from Fig.
-3, MolPIF (w/o mask) exhibited substructure distributions more similar to the reference molecules compared to the standard MolPIF.
+4, MolPIF (w/o mask) exhibited substructure distributions more similar to the reference molecules compared to the standard MolPIF.
 However, with the Laplace prior, the mask module showed selective improvements - it only benefited certain conventional properties (e.g., Vina Score, Vina Min, $\text{JS}_{\text{BL}}$, and $\text{JS}_{\text{BA}}$) while dramatically enhancing substructure generation.
@@
 This capability enabled arbitrary specification of fixed atoms within given molecules while allowing the model to generate desired substituent regions.
-Figure 4 presents selected lead optimization cases performed on reference molecules targeting 1umd, 3ZCW, and 6KZZ.
+Fig.
+5 presents selected lead optimization cases performed on reference molecules targeting 1umd, 3ZCW, and 6KZZ.
 The target 1umd was selected from the CrossDocked test set, with fixed ligand atom selection following the protocol established by CBGBench [54].
-Figure 4a displays some molecular structures generated by MolPIF for pocket 1umd during lead optimization.
+Fig.
+5a displays some molecular structures generated by MolPIF for pocket 1umd during lead optimization.
 The reference ligand was partitioned into distinct regions for four subtasks: fragment growth, linker design, scaffold hopping, and side-chain decoration.
@@
 coli DNA gyrase B [67] (PDB ID: 6KZZ) (Fig.
-4b-d).
+5b-d).
 For Kinesin Eg5, we adopted the same side-chain fixation strategy as Delete [68] to maintain consistency.
@@
 Comprehensive evaluations demonstrate the framework’s effectiveness, showing significant improvements in both general molecular properties and local geometries compared to state-of-the-art baselines.
+Chemical space analysis confirms MoIPIF’s ability to accurately replicate and expand molecular structural diversity.
 Case studies on targets including 2v3r, 1l3l and HAT1 (6VO5) further validate MoIPIF’s ability to generate novel molecules with enhanced binding properties over reference compounds.
@@
 2:Sample $i\sim\mathcal{U}\{0,n-1\}$
-3:$t\Leftarrow i/n$
-4:$\boldsymbol{\theta}(\textbf{x}_{\text{data}})\in\{\textbf{a}\mid p(\textbf{x}|\textbf{a})=\delta(\textbf{x}-\textbf{x}_{\text{data}})\}$
-5:$f(t)\Leftarrow 1-\gamma^{t}$
-6:$\boldsymbol{\theta}_{t}\Leftarrow f(t)\boldsymbol{\theta}(\textbf{x}_{\text{data}})+(1-f(t))\boldsymbol{\theta}_{\text{prior}}$
+[…]
+[…]
+[…]
+[…]
 7:Sample $\textbf{m}\sim p(\textbf{x}|\boldsymbol{\theta}_{t})$
-8:$\hat{\boldsymbol{\theta}}\Leftarrow\boldsymbol{\Phi}(\textbf{m})$
-9:$\Delta t\Leftarrow 1/n$
-10:$\boldsymbol{\theta}_{t+\Delta t}\Leftarrow f(t+\Delta t)\boldsymbol{\theta}(\textbf{x}_{\text{data}})+(1-f(t+\Delta t))\boldsymbol{\theta}_{\text{prior}}$
-11:$\hat{\boldsymbol{\theta}}_{t+\Delta t}\Leftarrow f(t+\Delta t)\hat{\boldsymbol{\theta}}+(1-f(t+\Delta t))\boldsymbol{\theta}_{\text{prior}}$
 […]
-13:$\boldsymbol{\Phi}\Leftarrow\boldsymbol{\Phi}-\alpha\nabla_{\boldsymbol{\Phi}}L_{t}$
+[…]
+[…]
+[…]
+[…]
+[…]
 Algorithm 2 Sampling procedure of PIF
 1:probability distribution $p(\textbf{x}|\boldsymbol{\theta})\in\mathcal{P}$, number of steps $n\in\mathbb{N}$, $\gamma\in\mathbb{R}^{+}$, $\boldsymbol{\theta}_{\text{prior}}\in\boldsymbol{\Theta}$, trained neural network $\boldsymbol{\Phi}$
-2:$\boldsymbol{\theta}_{0}\Leftarrow\boldsymbol{\theta}_{\text{prior}}$
+[…]
 3:for $i=0$ to $n-1$ do
@@
 As shown in Fig.
-5, our model generated samples that aligned well with the original distributions, exhibiting fewer outliers and reasonable coverage compared to baseline methods.
+6, our model generated samples that aligned well with the original distributions, exhibiting fewer outliers and reasonable coverage compared to baseline methods.
 These results suggested our approach had the ability to effectively handle both continuous and discrete patterns.